You purchase 26 call option contracts with a strike price of $140 and a premium of $4.35. Assume the stock price at expiration is $152.00.

a. What is your dollar profit? (Do not round intermediate calculations.)

b. What is your dollar profit if the stock price is $137.95? (A negative value should be indicated by a minus sign. Do not round intermediate calculations.)
Answer #1

(a) Value of call option at expiry = max (0, spot at expiry - strike price)

Value of call option at expiry = max (0, 152 - 140) = max (0,12) = 12

Profit per option = value of option at expiry - premium paid

Profit per option = 12 - 4.35 = $ 7.65

Dollar profit = profit per option x number of options

Dollar profit = 7.65 x 2600 = $ 19,890

(Note that if nothing is mentioned, each option contract is assumed to be of 100 options)

(b) Value of call option at expiry = max (0, spot at expiry - strike price)

Value of call option at expiry = max (0, 137.95 - 140) = max (0,-2.05) = 0

Profit / (Loss) per option = value of option at expiry - premium paid

Profit / (Loss) per option = 0 - 4.35 = ($4.85)

Dollar loss = loss per option x number of options

Dollar loss= -4.35 x 2600 = (-$ 11,310)

(Note that if nothing is mentioned, each option contract is assumed to be of 100 options)

Thumbs up please if satisfied. :)

Comment for more doubts in above question